OpenStudy (camerondoherty):

Solve the following system: y = x + 3 4x + y = 18 (7, 4) (-4, 7) (4, 7) (4, -7)

jimthompson5910 (jim_thompson5910):

Start with the second equation 4x + y = 18 and plug in y = x+3 4x + y = 18 4x + x+3 = 18 ... replace y with x+3, solve for x 5x+3 = 18 ... ... ... x = ???

OpenStudy (camerondoherty):

3

jimthompson5910 (jim_thompson5910):

Then plug that into y = x+3

jimthompson5910 (jim_thompson5910):

y = x+3 y = 3+3 ... plug in x = 3 (ie replace x with 3) y = 6

OpenStudy (camerondoherty):

y=6?

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (camerondoherty):

oh ok Thank You!

jimthompson5910 (jim_thompson5910):

So the solution is the ordered pair (3,6)

OpenStudy (camerondoherty):

but thts not one of the options...

jimthompson5910 (jim_thompson5910):

and there's a typo here since that's not a choice but I'm not surprised since I see this sort of thing a lot with online problems

OpenStudy (camerondoherty):

maybe u add 1 to each?

jimthompson5910 (jim_thompson5910):

double check to make sure all the numbers have been copied correctly

OpenStudy (camerondoherty):

ok hold on

OpenStudy (camerondoherty):

yea its copied correctly :/

OpenStudy (camerondoherty):

http://prntscr.com/3yt0e3

jimthompson5910 (jim_thompson5910):

then it's a typo on their part (your teacher or whoever wrote this problem)

jimthompson5910 (jim_thompson5910):

and there's not much you can do other than contact the teacher about it

OpenStudy (camerondoherty):

ok thank you ill do tht

OpenStudy (camerondoherty):

Thx Again!

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (camerondoherty):

ugh my teachers on vacation for the 4th of july until Sunday -_-

jimthompson5910 (jim_thompson5910):

no fair..you don't get the luxury of vacation

jimthompson5910 (jim_thompson5910):

oh well, hopefully you can still reach him/her on time to address the issue

OpenStudy (camerondoherty):

lol ill just put the one tht add 1 to each and if it marks it wrong ill call her later

OpenStudy (camerondoherty):

Thanks again for all your help @jim_thompson5910 i appreciate it :)

jimthompson5910 (jim_thompson5910):

ok I guess that works out and you're welcome